What is the growth rate for
P=3.2e^(0.03t)

Answer by stanbon(75887)

What is the growth rate for
P=3.2e^(0.03t)
---
Growth rate = e^0.03-1 = 1.03-1 = 3%
=====================================
Cheers,
Stan H.
====================